OpenStudy (anonymous):

To solve an equation of the form a b^x = d, you should first divide both sides by the coefficient _____.

OpenStudy (anonymous):

or you could just take the xth root of both sides and end up with B=xth root of D

OpenStudy (anonymous):

a is the coefficent in this problem.

OpenStudy (anonymous):

after dividing both sides by a, then you can take the xth root, or use properties of logarithms depending on which variable you are solving for

OpenStudy (anonymous):

divide by a then take the log

OpenStudy (anonymous):

for both sides

OpenStudy (anonymous):

B^x = d/a logB^x = log(d/a) xlogB = log(d/a) x= log(d/a)/logb